OpenStudy (anonymous):

F(x)=ln(x) g(x)=e^x+1. F(g(x)) must then be ln(e^x+1) right? Can you simplify this?

ganeshie8 (ganeshie8):

Hey!

ganeshie8 (ganeshie8):

\(\ln(t)\) is the inverse of \(e^t\), so it simplifies to : \[\ln (e^{x+1}) = x+1\]